# Beyond Noisy-TVs: Noise-Robust Exploration Via Learning Progress Monitoring

### v1

@@ Abstract
When there exists an unlearnable source of randomness (noisy-TV) in the environment, a naively intrinsic reward driven exploring agent gets stuck at that source of randomness and fails at exploration. Intrinsic reward based on uncertainty estimation or distribution similarity, while eventually escapes noisy-TVs as time unfolds, suffers from poor sample efficiency and high computational cost. Inspired by recent findings from neuroscience that humans monitor their improvements during exploration, we propose a novel method for intrinsically-motivated exploration, named Learning Progress Monitoring (LPM). During exploration, LPM rewards model improvements instead of prediction error or novelty, effectively rewards the agent for observing learnable transitions rather than the unlearnable transitions. We introduce a dual-network design that uses an error model to predict the expected prediction error of the dynamics model in its previous iteration, and use the difference between the model errors of the current iteration and previous iteration to guide exploration. We theoretically show that the intrinsic reward of LPM is zero-equivariant and a monotone indicator of Information Gain (IG), and that the error model is necessary to achieve monotonicity correspondence with IG. We empirically compared LPM against state-of-the-art baselines in noisy environments based on MNIST, 3D maze with 160x120 RGB inputs, and Atari. Results show that LPM’s intrinsic reward converges faster, explores more states in the maze experiment, and achieves higher extrinsic reward in Atari. This conceptually simple approach marks a shift-of-paradigm of noise-robust exploration. For code to reproduce our experiments, see https://github.com/Akuna23Matata/LPM_exploration
@@ 1 Introduction
While promising, this class of methods suffers from a long-standing challenge known as the noisy-TV problem (Burda et al., 2018). For example, a naively curious agent may become fixated on the unpredictability of static noise on a TV screen, wasting time gathering observations that do not improve its world model. Despite being a well-known issue (Schmidhuber, 1991; Burda et al., 2019), robust exploration in the presence of noisy-TVs remains an open problem. Existing works Mavor-Parker et al. (2022); Wang et al. (2024); Jiang et al. (2025) focused on isolating the uncertainty that can be reduced by more data (known as epistemic uncertainty) from the uncertainty that is irreducible by more data (aleatoric uncertainty), and filter the intrinsic reward signals to only encourage exploring the former type of uncertainty, which is also referred as novelty. However, distinguishing these two types of uncertainty is extremely challenging. Most approaches require either a strong prior over the state space (Wang et al., 2024; Jiang et al., 2025; Henaff et al., 2023; Zhang et al., 2021) or massive amounts of data (Burda et al., 2018; Mavor-Parker et al., 2022) before the agent can reliably filter out noise. This slow convergence causes early exploration to be dominated by noise, wasting large numbers of samples and limiting the agent’s ability to explore effectively in complex, real-world environments.
Recent neuroscience findings Ten et al. (2021) demonstrate that humans naturally monitor learning progress during exploration. During exploration, humans tend to observe the transitions that makes them learn the most about the dynamics. Because observing an unlearnable transition does not produce learning progress, this strategy is naturally robust to noisy-TV distractions and promotes efficient exploration by directing effort toward the most informative experiences. In the context of intrinsic reward, this insight suggests that we should reward model improvement rather than prediction error or state novelty. Inspired by this study, we propose a fundamentally different approach to intrinsic motivation, called Learning Progress Monitoring (LPM). In this work, we first develop a formal definition of the noisy-TV problem, then introduce our solution and complete algorithm, LPM, as a new form of intrinsic reward. We then provide a theoretical analysis showing that our approach is monotonically related to information gain. Finally, we conduct comprehensive experiments to evaluate the empirical performance of LPM and demonstrate its advantages over existing intrinsic motivation methods. We summarize our contributions as the followings:
@@ 3 Learning Progress Monitoring
Subsequently, we consider a model-based reinforcement learning setting, where the agent periodically updates a dynamics model $f_{\theta}$ that gives the prediction of next observation $\hat{o}_{t+1}$ given the current observation $o_{t}$ and action $a_{t}$: $\hat{o}_{t+1}=f^{(\tau)}_{\theta}(o_{t},a_{t})$ Let $\mathcal{B}$ consists of $(o_{t},a_{t},o_{t+1})$ be the replay buffer of transition dynamics used to fit $f_{\theta}$. Upon observing the true next observation $o_{t+1}$, we can compute the difference between the true and predicted next observations. We define the log Mean Squared Error (MSE) of $\tau$-th dynamics model’s prediction at time step $t$ to be $\varepsilon_{t}^{(\tau)}$:
@@ 5 Experiment
Ensemble (Lakshminarayanan et al., 2017) $\mathrm{Var}_{i}[f_{i}(s_{t},a_{t})]$ Curiosity
@@ 5.1 Noisy MNIST
To answer the first research question, we evaluate our method using the noisy MNIST dataset following the experimental setup from (Mavor-Parker et al., 2022; Pathak et al., 2019; Mazzaglia et al., 2022).The noisy MNIST (Deng, 2012) environment contains separated deterministic and stochastic state transitions, allowing us to observe the behavior characteristic across different intrinsic motivation methods. The deterministic state starts with initial state of MNIST image of 0, and transit to next state that is exactly the same as initial state. While the stochastic state starts with initial state of MNIST image 1, and transit to randomly sampled next state of MNIST image 2-9. We focus on two representative baselines: AMA, a noise-robust curiosity-driven method, and Episodic Novelty Through Temporal Distance (EDT) Jiang et al. (2025), an similarity based episodic bonus method. These two baselines serve as representatives of their respective families, providing a clear comparison of how different intrinsic motivation mechanisms behave as more transitions are observed.
@@ 5.2 MiniWorld
To answer the second research question—whether LPM explores more novel states in pure exploration tasks compared with existing methods—we evaluate our method in the 3D MiniWorld environment Chevalier-Boisvert et al. (2023) with high-dimensional observations (160×120 RGB). This experiment focuses on pure exploration performance, measuring the agent’s ability to discover and visit diverse states when guided solely by intrinsic motivation methods. We use a 3-room N-shaped MiniWorld maze with 2,176 visitable states. Each room has distinct textures to increase visual diversity, making exploration visually complex and closer to real-world environments. Additionally, we test exploration consistency across different noise levels, challenging existing methods that may fail in realistic visual environments where unpredictable stimuli can trigger noisy-TV behavior by creating three experimental conditions to evaluate robustness: (1) a deterministic setting with fixed textures and transitions, (2) a state noise setting where the middle room exits a noisy wall displaying random pixel noise at each timestep to test whether agents fixate on unpredictable visual stimuli, and (3) an action noise setting where selecting the idle action replaces the true observation with a random CIFAR-10 image Krizhevsky et al. (2009), simulating action-triggered noise. This follows similar noise setting as MiniGrid environment from Mavor-Parker et al. (2022).
@@ 5.3 Atari
To answer the third research question, we evaluate our method on several hard exploration Atari games identified in prior work Wang et al. (2024); Bellemare et al. (2016); Wang et al. (2023), including Space Invader and Ms PacMan. To systematically test noise robustness while preserving meaningful gameplay, we extend each game’s action space by adding two idle (no-op) actions. This preserves the utility of idle actions while allowing for controlled noise injection. For each game, we construct two environment variants: (1) the original Atari game without modifications, and (2) an action noise variant, where selecting an idle action replaces the true observation with a random CIFAR-10 image Krizhevsky et al. (2009), following the noise-injection setup from Mavor-Parker et al. (2022). This design allows us to directly evaluate how intrinsic motivation methods handle noise when exploring challenging, goal-directed tasks.
@@ 5.4 Computational overhead
While LPM employs a dual-network structure to estimate learning progress, its computational complexity remains modest when compared to baseline methods such as AMA Mavor-Parker et al. (2022), Ensemble Lakshminarayanan et al. (2017), and EME Wang et al. (2024). AMA uses a double-headed neural network, which require similar computational cost with LPM. Ensemble and EME both require multiple models and hence require significantly higher computational costs.
Figure 4: Average and standard deviation extrinsic rewards achieved by each method across deterministic and stochastic Atari environments. Each entry represents the aggregated result from 128 random seeds.
@@ 6.2 Exploration through Episodic Bonus
Episodic bonus methods provide intrinsic motivation by rewarding state novelty within the current episode (Zhang et al., 2021; Henaff et al., 2023). These approaches generally fall into two categories: Count-based methods (Raileanu & Rocktäschel, 2020; Flet-Berliac et al., 2021; Mu et al., 2022; Zhang et al., 2021) generate a positive intrinsic reward signal whenever the agent visits a previously unseen state within an episode, thereby encouraging exploration of novel states. In contrast, similarity-based methods (Savinov et al., 2018; Wang et al., 2024; Jiang et al., 2025; Yang et al., 2024) define novelty as a continuous function of the distance between the current state and past visited states, using metrics such as temporal distance (Jiang et al., 2025), Kullback–Leibler divergence (Wang et al., 2024), or Euclidean distance (Henaff et al., 2022).
@@ 6.3 Information-Theoretic Exploration
Information-theoretic exploration methods reward information gain, encouraging the agent to take actions that maximally reduce its uncertainty about the environment dynamics (Houthooft et al., 2016). This is typically formalized as information gain (IG), the KL divergence between the agent’s posterior and prior beliefs upon receiving a new observation. By directly maximizing IG, these approaches have strong theoretical guarantees that the collected data will be maximally informative for learning state transitions. In practice, estimating information gain requires maintaining a posterior distribution over the environment, which is typically achieved using Gaussian processes (GPs) (Bottero et al., 2022; Sui et al., 2015; Hennig & Schuler, 2012; Hernández-Lobato et al., 2014) or Bayesian neural networks (BNNs) (Houthooft et al., 2016; Mazzaglia et al., 2022; Blau et al., 2019; Li & Faisal, 2021). While theoretically appealing, these methods are computationally expensive and hard to scale to high-dimensional, visual domains such as Atari or robotics.
@@ 7 Conclusion
Motivated by recent neuroscience findings Ten et al. (2021), we propose using learning progress as intrinsic reward to motivate reinforcement learning exploration. We designed the Learning Progress Monitoring (LPM) method with a dual network structure to monitor learning progress. We then thoroughly tested our method in environments with and without extrinsic rewards, state-based noise, and action-based noise. Our results demonstrate that learning progress as intrinsic reward, compared to existing intrinsic motivation methods that use prediction error or distribution similarity, achieves higher exploration efficiency while being naturally robust to noise. This makes our LPM method particularly suitable for realistic environments where deterministic and stochastic elements coexist.
@@
We provide full implementation details in Appendix B, including network architectures, hyperparameters, environment configurations, and noise injection procedures. All random seeds used in our experiments are specified.
@@ A.2 MLE satisfies the $\theta_{D}$ condition
Appendix B Implementation Details
@@
B.1 Noisy MNIST Implementation Details

### v2

@@ Abstract
When there exists an unlearnable source of randomness (noisy-TV) in the environment, a naively intrinsic reward driven exploring agent gets stuck at that source of randomness and fails at exploration. Intrinsic reward based on uncertainty estimation or distribution similarity, while eventually escapes noisy-TVs as time unfolds, suffers from poor sample efficiency and high computational cost. Inspired by recent findings from neuroscience that humans monitor their improvements during exploration, we propose a novel method for intrinsically-motivated exploration, named Learning Progress Monitoring (LPM). During exploration, LPM rewards model improvements instead of prediction error or novelty, effectively rewards the agent for observing learnable transitions rather than the unlearnable transitions. We introduce a dual-network design that uses an error model to predict the expected prediction error of the dynamics model in its previous iteration, and use the difference between the model errors of the current iteration and previous iteration to guide exploration. We theoretically show that the intrinsic reward of LPM is zero-equivariant and a monotone indicator of Information Gain (IG), and that the error model is necessary to achieve monotonicity correspondence with IG. We empirically compared LPM against state-of-the-art baselines in noisy environments based on MNIST, 3D maze with 160x120 RGB inputs, and Atari. Results show that LPM’s intrinsic reward converges faster, explores more states in the maze experiment, and achieves higher extrinsic reward in Atari. This conceptually simple approach marks a shift-of-paradigm of noise-robust exploration.
footnotetext: Code available at https://github.com/Akuna23Matata/LPM_exploration
@@ 1 Introduction
While promising, this class of methods suffers from a long-standing challenge known as the noisy-TV problem (Burda et al., 2018). For example, a naively curious agent may become fixated on the unpredictability of static noise on a TV screen, wasting time gathering observations that do not improve its world model. Despite being a well-known issue (Schmidhuber, 1991; Burda et al., 2019), robust exploration in the presence of noisy-TVs remains an open problem. Existing works (Mavor-Parker et al., 2022; Wang et al., 2024; Jiang et al., 2025) focused on isolating the uncertainty that can be reduced by more data (known as epistemic uncertainty) from the uncertainty that is irreducible by more data (aleatoric uncertainty), and filter the intrinsic reward signals to only encourage exploring the former type of uncertainty, which is also referred as novelty. However, distinguishing these two types of uncertainty is extremely challenging. Most approaches require either a strong prior over the state space (Wang et al., 2024; Jiang et al., 2025; Henaff et al., 2023; Zhang et al., 2021) or massive amounts of data (Burda et al., 2018; Mavor-Parker et al., 2022) before the agent can reliably filter out noise. This slow convergence causes early exploration to be dominated by noise, wasting large numbers of samples and limiting the agent’s ability to explore effectively in complex, real-world environments.
Recent neuroscience findings (Ten et al., 2021) demonstrate that humans naturally monitor learning progress during exploration. During exploration, humans tend to observe the transitions that makes them learn the most about the dynamics. Because observing an unlearnable transition does not produce learning progress, this strategy is naturally robust to noisy-TV distractions and promotes efficient exploration by directing effort toward the most informative experiences. In the context of intrinsic reward, this insight suggests that we should reward model improvement rather than prediction error or state novelty. Inspired by this study, we propose a fundamentally different approach to intrinsic motivation, called Learning Progress Monitoring (LPM). In this work, we first develop a formal definition of the noisy-TV problem, then introduce our solution and complete algorithm, LPM, as a new form of intrinsic reward. We then provide a theoretical analysis showing that our approach is monotonically related to information gain. Finally, we conduct comprehensive experiments to evaluate the empirical performance of LPM and demonstrate its advantages over existing intrinsic motivation methods. We summarize our contributions as the followings:
@@ 3 Learning Progress Monitoring
Subsequently, we consider a model-based reinforcement learning setting, where the agent periodically updates a dynamics model $f_{\theta}$ that gives the prediction of next observation $\hat{o}_{t+1}$ given the current observation $o_{t}$ and action $a_{t}$: $\hat{o}_{t+1}=f^{(\tau)}_{\theta}(o_{t},a_{t})$. Let $\mathcal{B}$ consists of $(o_{t},a_{t},o_{t+1})$ be the replay buffer of transition dynamics used to fit $f_{\theta}$. Upon observing the true next observation $o_{t+1}$, we can compute the difference between the true and predicted next observations. We define the log Mean Squared Error (MSE) of $\tau$-th dynamics model’s prediction at time step $t$ to be $\varepsilon_{t}^{(\tau)}$:
@@ 5 Experiment
Ensemble (Pathak et al., 2019) $\mathrm{Var}_{i}[f_{i}(s_{t},a_{t})]$ Curiosity
@@ 5.1 Noisy MNIST
To answer the first research question, we evaluate our method using the noisy MNIST dataset following the experimental setup from (Mavor-Parker et al., 2022; Pathak et al., 2019; Mazzaglia et al., 2022).The noisy MNIST (Deng, 2012) environment contains separated deterministic and stochastic state transitions, allowing us to observe the behavior characteristic across different intrinsic motivation methods. The deterministic state starts with initial state of MNIST image of 0, and transit to next state that is exactly the same as initial state. While the stochastic state starts with initial state of MNIST image 1, and transit to randomly sampled next state of MNIST image 2-9. We focus on two representative baselines: AMA, a noise-robust curiosity-driven method, and Episodic Novelty Through Temporal Distance (EDT) (Jiang et al., 2025), an similarity based episodic bonus method. These two baselines serve as representatives of their respective families, providing a clear comparison of how different intrinsic motivation mechanisms behave as more transitions are observed.
@@ 5.2 MiniWorld
To answer the second research question—whether LPM explores more novel states in pure exploration tasks compared with existing methods—we evaluate our method in the 3D MiniWorld environment (Chevalier-Boisvert et al., 2023) with high-dimensional observations (160×120 RGB). This experiment focuses on pure exploration performance, measuring the agent’s ability to discover and visit diverse states when guided solely by intrinsic motivation methods. We use a 3-room N-shaped MiniWorld maze with 2,176 visitable states. Each room has distinct textures to increase visual diversity, making exploration visually complex and closer to real-world environments. Additionally, we test exploration consistency across different noise levels, challenging existing methods that may fail in realistic visual environments where unpredictable stimuli can trigger noisy-TV behavior by creating three experimental conditions to evaluate robustness: (1) a deterministic setting with fixed textures and transitions, (2) a state noise setting where the middle room exits a noisy wall displaying random pixel noise at each timestep to test whether agents fixate on unpredictable visual stimuli, and (3) an action noise setting where selecting the idle action replaces the true observation with a random CIFAR-10 image (Krizhevsky et al., 2009), simulating action-triggered noise. This follows similar noise setting as MiniGrid environment from (Mavor-Parker et al., 2022).
@@
5.3 Mountain Car
To complement the MiniWorld experiment—which evaluates state coverage in a discrete action setting, we further conducted a continuous control exploration task. We constructed two variants of the MountainCar-Continuous environment to evaluate LPM under sparse rewards, continuous environments.
The deterministic environment includes the standard two-dimensional state (position, velocity) and one-dimensional continuous action. We remove all original rewards and place three reward points at random locations per run, creating a sparse-reward continuous control task. The stochastic environment extend the action space with an additional noisy action dimension. When this extended action is chosen with a value greater than zero, the environment freezes the true state and outputs a random observation sampled uniformly from [-1,1]. During testing, we separate the MountainCar state space into 100 bins and measure how many unique bins each method visits during training with 5 random seeds. As shown in Table 2, LPM can motivate exploration in deterministic continues environment, attains the best performance under stochasticity and shows the smallest degradation, with only a 12.4% drop, which is far lower than all baselines. This highlights LPM’s strong robustness to stochasticity.
Table 2: State coverage comparison in the MountainCar continuous-control environment with deterministic and stochastic settings. Best performance in each task is in bold.
Method Deterministic Coverage (%) Stochastic Coverage (%) Drop (%)
LPM $76.50\pm 9.08$ $\mathbf{67.04\pm 14.60}$ $\mathbf{12.4}$
Ensemble $\mathbf{91.22\pm 2.04}$ $61.02\pm 5.03$ $33.1$
RND $45.50\pm 14.53$ $28.00\pm 10.10$ $38.5$
AMA $33.00\pm 12.31$ $13.20\pm 4.62$ $60.0$
EDT $82.16\pm 13.57$ $53.52\pm 10.53$ $34.9$
EME $89.16\pm 3.40$ $32.46\pm 11.31$ $63.6$
IDF $90.92\pm 5.13$ $12.80\pm 3.19$ $85.9$
@@ 5.4 Atari
To answer the third research question, we evaluate our method on several hard exploration Atari games identified in prior work (Wang et al., 2024; Bellemare et al., 2016; Wang et al., 2023), including Space Invader and Ms PacMan. To systematically test noise robustness while preserving meaningful gameplay, we extend each game’s action space by adding two idle (no-op) actions. This preserves the utility of idle actions while allowing for controlled noise injection. For each game, we construct two environment variants: (1) the original Atari game without modifications, and (2) an action noise variant, where selecting an idle action replaces the true observation with a random CIFAR-10 image (Krizhevsky et al., 2009), following the noise-injection setup from (Mavor-Parker et al., 2022). This design allows us to directly evaluate how intrinsic motivation methods handle noise when exploring challenging, goal-directed tasks.
@@
5.5 Montezuma’s Revenge
Beyond the standard Atari benchmarks, we further evaluate LPM on Montezuma’s Revenge, one of the most challenging exploration games where most intrinsic-motivation methods fail to obtain any score. While prior experiments showed that LPM is highly robust to environmental stochasticity, this test examines whether LPM can still guide exploration in extremely sparse-reward settings. Remarkably in Figure 5, LPM is able to generate meaningful exploration behavior and achieves non-trivial performance within 20M exploration steps, whereas RND requires 50M exploration steps. Since NGU(Badia et al., 2020) has no official implementation, we re-implemented it following the original paper. NGU requires days of wall-clock training, so we report only 50M steps, during which it failed to obtain any extrinsic reward. This result suggests that LPM is not only noise-robust but also capable of driving exploration in hard long-horizon tasks such as Montezuma’s Revenge.
@@ 5.6 Computational overhead
While LPM employs a dual-network structure to estimate learning progress, its computational complexity remains modest when compared to baseline methods such as AMA (Mavor-Parker et al., 2022), Ensemble (Pathak et al., 2019), and EME (Wang et al., 2024). AMA uses a double-headed neural network, which require similar computational cost with LPM. Ensemble and EME both require multiple models and hence require significantly higher computational costs.
Figure 4: Average and standard deviation extrinsic rewards achieved by each method across deterministic and stochastic Atari environments with 128 random seeds.
@@ 6.2 Exploration through Episodic Bonus
Episodic bonus methods provide intrinsic motivation by rewarding state novelty within the current episode (Zhang et al., 2021; Henaff et al., 2023). These approaches generally fall into two categories: Count-based methods (Raileanu & Rocktäschel, 2020; Flet-Berliac et al., 2021; Mu et al., 2022; Zhang et al., 2021; Kapturowski et al., 2024) generate a positive intrinsic reward signal whenever the agent visits a previously unseen state within an episode, thereby encouraging exploration of novel states. In contrast, similarity-based methods (Savinov et al., 2018; Wang et al., 2024; Jiang et al., 2025; Yang et al., 2024) define novelty as a continuous function of the distance between the current state and past visited states, using metrics such as temporal distance (Jiang et al., 2025), Kullback–Leibler divergence (Wang et al., 2024), or Euclidean distance (Henaff et al., 2022).
@@ 6.3 Information-Theoretic Exploration
Information-theoretic exploration methods reward information gain, encouraging the agent to take actions that maximally reduce its uncertainty about the environment dynamics (Houthooft et al., 2016; Rhinehart et al., 2021). This is typically formalized as information gain (IG), the KL divergence between the agent’s posterior and prior beliefs upon receiving a new observation. By directly maximizing IG, these approaches have strong theoretical guarantees that the collected data will be maximally informative for learning state transitions. In practice, estimating information gain requires maintaining a posterior distribution over the environment, which is typically achieved using Gaussian processes (GPs) (Bottero et al., 2022; Sui et al., 2015; Hennig & Schuler, 2012; Hernández-Lobato et al., 2014) or Bayesian neural networks (BNNs) (Houthooft et al., 2016; Mazzaglia et al., 2022; Blau et al., 2019; Li & Faisal, 2021). While theoretically appealing, these methods are computationally expensive and hard to scale to high-dimensional, visual domains such as Atari or robotics.
Figure 5: Exploration performance on Montezuma’s Revenge. LPM enables the agent to achieve meaningful extrinsic reward within 20M steps, whereas baselines require 40M+ steps to reach comparable scores. Center: LPM’s intrinsic reward gradually decays as exploration saturates, mirroring its behavior on Noisy MNIST. Once the agent enters a previously unseen region around 20M steps, the intrinsic reward spikes, guiding the agent toward high-reward states. Right: LPM steadily expands its state coverage and discovers novel states demonstrating its ability to drive exploration even in extremely sparse-reward environments.
@@ 7 Conclusion
Motivated by recent neuroscience findings (Ten et al., 2021), we propose using learning progress as intrinsic reward to motivate reinforcement learning exploration. We designed the Learning Progress Monitoring (LPM) method with a dual network structure to monitor learning progress. We then thoroughly tested our method in environments with and without extrinsic rewards, state-based noise, and action-based noise. Our results demonstrate that learning progress as intrinsic reward, compared to existing intrinsic motivation methods that use prediction error or distribution similarity, achieves higher exploration efficiency while being naturally robust to noise. This makes our LPM method particularly suitable for realistic environments where deterministic and stochastic elements coexist.
Ethics Statement
All authors have read and adhere to the ICLR Code of Ethics.
@@
Detailed experimental setups, hyperparameters, architectural specifications, and links to our source code are documented in Appendix C.
Acknowledgment
This work was supported in part by the NSF Grant #2239458 and an UC Merced Fall 2023 Climate Action Seed Competition Grant. Any opinions, findings, and conclusions expressed in this material are those of the authors and do not necessarily reflect the views of the funding agencies.
@@ A.2 MLE satisfies the $\theta_{D}$ condition
Appendix B MiniGrid Experiment
We further evaluate LPM within the MiniGrid(Chevalier-Boisvert et al., 2023) framework, using the Lava Crossing environment. This maze-like domain serves as an ideal testbed for intrinsic motivation: while agents must explore to reach the goal, excessive or unregulated curiosity can be detrimental, driving the agent into fatal lava grids. As illustrated in Figure 6, both ICM and our proposed LPM demonstrate comparable performance, significantly outperforming the remaining baselines.
Figure 6: PPO Agent success rate across 5 random seeds.
Appendix C Implementation Details
@@
C.1 Noisy MNIST Implementation Details